OpenStudy (anonymous):

what is the product of sqrt t (6 - 2 sqrt 7)? A) 6 sqrt t - 2 sqrt 7 B) 6 sqrt t - 2 sqrt 7t C) 4t sqrt 7 D) 4 sqrt 7t

OpenStudy (zehanz):

\(\sqrt{t}(6-2\sqrt{7})\) is just like a(b-c). You can use the distributive property: a(b-c)=ab-ac.

OpenStudy (zehanz):

Also, use this rule for roots: \(\sqrt{a} \cdot \sqrt{b}=\sqrt{ab}\)

OpenStudy (anonymous):

but how would I use that with the equation? I don't understand :/

OpenStudy (zehanz):

OK, let's see what to do: \[\sqrt{t}(6-2\sqrt{7})=\sqrt{t} \cdot 6 - \sqrt{t}\cdot 2\sqrt{7}\] I used the distributive property (distributed √t between the two numbers in brackets).

OpenStudy (zehanz):

Tidy up a bit: \(6\sqrt{t}-2\sqrt{t}\sqrt{7}=6\sqrt{t}-2\sqrt{7t}\) Now I used the rule I mentioned about roots.

OpenStudy (anonymous):

Oh, okay Thank you @ZeHanz